OpenStudy (anonymous):

f(x)=lnsqrt(4x-5/9x+9) find the deriv of f

OpenStudy (lgbasallote):

chain rule will tell you... let y = ln(sqrt(4x-5/9x+9)) x = sqrt(4x-5/9x+9) w = 4x -5 z = 9x + 9 so (zdw - wdz)/xz^2 just substitute back

OpenStudy (lgbasallote):

btw...dw = 4 and dz = 9